# Fermi-LAT Discovery of a Gamma-ray Outburst from the Peculiar Compact Steep Spectrum Radiogalaxy 3C 216

### v1

@@ 3 Neil Gehrels Swift XRT and UVOT data
Following the powerful gamma-ray outburst of 3C 216, detected on 2023-05-01, (Giacchino et al., 2023), a Target of Opportunity (ToO) request was sent to the Swift Gamma-Ray Burst Mission (Gehrels et al., 2004). Swift executed five visits of the target, on 2023-05-03, 04, 06, 08 and 09, accumulating observations with both the XRT and UVOT instruments. The observations were executed with a roughly regular spacing of $1.5$ days between subsequent visits. During the observations, the XRT instrument worked in photon counting mode, while UVOT performed a sequence of exposures using the $V$, $B$, $U$, $UW1$, $UM2$ and $UW2$ photometric filters.
@@ 3.1 Analysis of Swift data
3C 216 was detected in X-ray/UV/optical frequencies with values range of the X-ray flux of $\Phi_{X}\sim(2.3-3.2)\times 10^{-12}\,\mathrm{erg/cm^{2}/s}$, UV magnitudes in the range $14.6$-$15.8$, and optical ones in the range $15.2$-$16.5$. In Table 3 we report also the fluxes in different filters for UVOT, and the flux and photon index for XRT. We report also the results for archival observations carried out on 2010-10-21.
Time 2010-11-21 2023-05-03b 2023-05-04b 2023-05-06b 2023-05-08c 2023/05/09c
UVOT UVV $(5.69\pm 0.82)^{a}$ $(1.29\pm 0.02)^{b}$ $(1.20\pm 0.02)^{b}$ $(0.92\pm 0.02)^{b}$ $(6.00\pm 0.15)^{c}$ $(4.31\pm 0.11)^{c}$
UVOT UBB $(7.69\pm 0.70)^{a}$ $(1.46\pm 0.03)^{b}$ $(1.47\pm 0.03)^{b}$ $(1.10\pm 0.03)^{b}$ $(6.84\pm 0.22)^{c}$ $(4.97\pm 0.16)^{c}$
UVOT UUU $(4.92\pm 0.41)^{a}$ $(1.48\pm 0.02)^{b}$ $(1.41\pm 0.03)^{b}$ $(1.12\pm 0.02)^{b}$ $(6.70\pm 0.22)^{c}$ $(5.27\pm 0.17)^{c}$
UVOT UM2 $(1.35\pm 0.34)^{a}$ $(1.65\pm 0.03)^{b}$ $(1.60\pm 0.03)^{b}$ $(1.33\pm 0.03)^{b}$ $(8.87\pm 0.26)^{c}$ $(6.31\pm 0.16)^{c}$
UVOT UW1 $(2.38\pm 0.40)^{a}$ $(1.75\pm 0.03)^{b}$ $(1.74\pm 0.03)^{b}$ $(1.35\pm 0.03)^{b}$ $(9.45\pm 0.28)^{c}$ $(7.18\pm 0.19)^{c}$
UVOT UW2 $(2.16\pm 0.38)^{a}$ $(1.64\pm 0.04)^{b}$ $(1.57\pm 0.05)^{b}$ $(1.43\pm 0.04)^{b}$ $(9.29\pm 0.40)^{c}$ $(7.21\pm 0.29)^{c}$
XRT $(1.54\pm 0.41)^{c}$ $(3.24\pm 0.42)^{c}$ $(2.36\pm 0.42)^{c}$ $(2.40\pm 0.49)^{c}$ $(2.26\pm 0.61)^{c}$ $(2.84\pm 0.72)^{c}$
$\Gamma_{X}$ $1.24\pm 0.57$ $1.91\pm 0.11$ $2.03\pm 0.16$ $1.58\pm 0.17$ $1.62\pm 0.23$ $1.29\pm 0.18$
Table 3: Table of fluxes observed by Swift for UVOT (in six bands) and XRT (from $0.3$ to $10$ keV). Photon index $\Gamma_{X}$ for spectral model in X-ray. The error are $1\sigma$. The fluxes are in erg/cm2/sec.a is ×$\times 10^{-14}$. b is ×$\times 10^{-11}$. c is ×$\times 10^{-12}$.
@@
The UVOT data were processed according to a standard UVOT software analysis, using heasoft-6.32.1444https://heasarc.gsfc.nasa.gov/docs/software/heasoft/. At first, we combined the exposures of each filter with uvotimsum, so to obtain one image per visit per filter. Then we extracted the flux and magnitude in each band-pass, together with their associated errors, using the task uvotdetect. The source was clearly detected in a high state, achieving a statistical significance larger than $20\sigma$ in all the visits and with all the photometric filters. The first Swift observation recorded a $U$-band magnitude of $(14.67\pm 0.02)\,$mag, more than 3 magnitudes brighter than the archival value of 17.8817.8817.88\,17.88mag reported in literature in the same band (Ryle & Sandage, 1964), while subsequent observations traced a steadily decreasing trend. In spite of the long time elapsed since this archival photometric determination, the source has also been observed more recently in ugriz photometry by the Sloan Digital Sky Survey, resulting in even fainter magnitude values in all pass-bands (Adelman-McCarthy et al., 2008), and by Swift itself in 2010, obtaining magnitudes of the order of $19$. We can therefore conclude that the UVOT magnitudes observed during the gamma-ray outburst correspond to a high state of the source. In order to recover the intrinsic state, we corrected the observed fluxes, accounting for the effects of foreground extinction due to the Milky Way’s interstellar medium. Since 3C 216 is located far away from the Galactic Plane, it is subject to a modest reddening effect, having $A_{V}=0.052$ mag, $A_{B}=0.069$ mag and $A_{U}=0.082$ mag (Schlafly & Douglas, 2011). To correct the $UV$ pass-bands, we derived the extinction coefficients for the $UW2$, $UM2$ and $UW1$ according to Roming et al. (2009) and we compared them with the corresponding values estimated by Yi et al. (2023), who claim to have updated photometric information. Since the two methods provided consistent estimates, with the largest difference being a 2% smaller intrinsic flux in the $UM2$ band according to Yi et al. (2023), with respect to Roming et al. (2009), we eventually used the corrections derived by Yi et al. (2023).
@@
The XRT data were processed using xrtpipeline v3.7.0555https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/xrt.html. For each observation, spectral analysis of the source was performed within a $35$ arcsec radius of target, with background emission subtracted from an annular region between an inner radius $r_{in}\sim 100$ arcsec and an outer radius $r_{out}\sim 400$ arcsec, centered on the position of 3C 216. The spectra were then binned to ensure a minimum of one count per bin and modeled using the XSPEC v12.13.1e (Arnaud et al., 1999) package, employing the Cash statistic for minimization Cash (1979). Fluxes were extracted in the soft ($[0.5-2]$ keV) and hard ($[2-10]$ keV) bands. The adopted model is a power-law modified by neutral Galactic absorption $TB_{abs}\times z_{pow}$.
@@
Figure 5: Light curves from UVOT data across six filters (top panel), XRT data in two energy ranges (medium panel), and Fermi-LAT daily binned data (bottom panel) over the same observation period (PF).
In Fig. 5 we present the light curves across different frequency bands for the PF period: the UVOT data for the six filters are shown in the top panel, the XRT data for the soft and hard bands in the middle panel, and the LAT data in the bottom panel. The observed trend indicates a stable decreasing evolution in the light curves across all spectral bands, with the exception of the hard X-ray range. This behavior is consistent with a scenario involving an initial simultaneous injection of energy into both the low frequency synchrotron component and the $\gamma$-ray IC component, followed by a radiative cooling phase. During this phase, while the overall power of the source decreases, the frequency of the IC scattered radiation also reduces, moving the high energy spectral hump into the hard X-ray regime. This effect can be clearly appreciated by looking at the evolution of the SED of 3C 216, plotted in Fig. 6, which shows that a relative increase of the hard-X ray flux agrees with the predictions of a SSC model, where a distribution of relativistic charged particles loses energy through radiative cooling (Tramacere et al., 2011).
@@
Figure 6: SED for the broadband emission of 3C 216, spanning from UV to gamma-ray, based on five datasets observed during the PF period. The blue line represents the SSC model of high-energy photon emission computed with JetSeT. The green dots indicate the data collected by UVOT, XRT, and Fermi-LAT, along with their respective uncertainties. Upper limits are reported when $TS\leq 10$. Light gray points in the background represent archival, non-simultaneous data from literature, specifically from the following sources: 3C, NRAO, Ohio (Dixon, 1970), FIRST at VLA, Green Bank GB6 and North surveys, combined NRAO and Parkes survey, NVSS, VLBA, Planck ERCSC, PCCS1F and PCCS2F catalogs, WMAP catalog, TWOMASS and WISE infrared catalogs, USNO and SDSS2,6,7, optical surveys, Einstein and ROSAT X-ray source catalogs, Swift XRT SWXRT1 and 1SXPS catalogs (D’Elia et al., 2013; Evans et al., 2014), Chandra ACIS source catalog (Massaro et al., 2015), and the Fermi-LAT DR3 source catalog (Abdollahi et al., 2022) data. These data were extracted from the SSDC SED Builder and NED archives.
4 Discussion and conclusions
Due to its overall steep radio spectrum and relatively compact morphology, 3C 216 is classified as a CSS source. In the youth scenario (Fanti et al., 1995), these sources owe their compactness to a young age and represent an early evolutionary stage towards a fully developed radio galaxy size. CSS sources, and their even more compact (and younger) siblings (GHz Peaked Spectrum, or GPS, sources, O’Dea et al. 1991), were predicted to be good candidates for GeV emission, as they combine on sub-galactic scales both the presence of recently injected relativistic particles and abundant photon fields from the central regions of their hosts (Stawarz et al., 2008). However, only a handful of CSS/GPS sources have been detected individually (Migliori et al., 2016; Principe et al., 2020) and even a stacking analysis has not revealed a collective signal from this population (Principe et al., 2021).
Except for a few outstanding sources detected because of their extreme proximity, a significant contribution from Doppler beamed components, such as relativistic jets, is required for CSS and GPS sources to be detected in gamma rays. This is suggested by elements such as the position in the luminosity-photon index diagram, the identification with quasar hosts, and the presence of variability. While Principe et al. (2021) already pointed out all these characteristics for 3C 216, the flare analyzed here reveals a level of flux variability significantly more extreme than anything observed before in this source.
For modeling and fitting radiative emission in 3C 216, we used the JetSeT tool (Tramacere, 2020). JetSeT is an open source C/Python framework for reproducing the radiative and acceleration processes acting in relativistic jets, and Galactic objects (beamed and unbeamed), allowing numerical models to be fit to the observed data. The included radiative processes in the tool are: synchrotron self-Compton, external Compton (EC) and EC against the CMB.
We decided to analyze the time variability of the multi-frequency SED to understand the physical process of the flare activity of 3C 216. We chose to construct our base-line model with the data set corresponding to 2023-05-03, because it is the earliest time for which we obtained simultaneous multi-frequency coverage of the source. We then tested the model on the following days, to check whether the SED evolution is consistent with an initial energy injection, followed by radiative losses and cooling. UVOT flux densities and X-ray data have been already explained in the previous section. In the gamma energy range, we considered the same Fermipy parameters of Table 1 for each day.
Since the SED evolves quite coherently in nearly all the observed pass-bands, instead of considering the possibility of external contribution to the IC component, we tested whether the status of the source could be reproduced by a SSC (Jones et al., 1974). In this case the seed photons for the IC process are the synchrotron photons produced by the same population of relativistic electrons that scatter them up to gamma-ray energies. We observed that the broad band shape of the SED could be explained by a leptonic plasma blob (the emitting region) with the characteristics reported in Table 4.
@@ 4 Discussion and conclusions
Table 4: Leptonic plasma blob characteristics.
The observed spectral shape of photons can be reproduced by a charged lepton population with a flat low-energy spectral index and a log-parabola energy distribution $f(\gamma)$. The required density, $N=1000\,\mathrm{cm^{-3}}$, is typical of the immediate environment surrounding an AGN, and the blob size is consistent with a variability timescale on the order of $1$ day in the observed frame. The number of emitting particles $N$ per unit volume is given by:
@@
where we adopted a constant value of $\gamma_{min}=2$ for the minimum Lorentz factor of the emitting particles energy distribution. The spectral law is defined by a log-parabola:
@@
A comparison of the adopted model with JetSeT (Massaro et al., 2006; Tramacere et al., 2009, 2011) and the broadband observed SED is illustrated in Fig. 6. Table 5 provides a summary of the changes in the spectral parameters required to account for the evolution of the SED.
Obs. Date $\gamma_{\mathrm{max}}$ $r$ $s$ $\gamma_{0}$
2023-05-03 $5\times 10^{5}$ $1.05$ $1$ 580
2023-05-04 $4\times 10^{5}$ $1.05$ $1.05$ $575$
2023-05-06 $4\times 10^{5}$ $1.07$ $1.07$ $540$
2023-05-08 $3\times 10^{5}$ $1.15$ $1.15$ $540$
2023-05-09 $3\times 10^{5}$ $1.15$ $1.2$ $530$
@@
Table 5: List of parameter model values from JetSeT tool that evolve over time. The reported parameters include time, the maximum energy of energy interval for the emitting electron distribution ($\gamma_{\mathrm{max}}$), the curvature of log-parabola ($r$), the spectral index of log-parabola ($s$), and the reference energy of log-parabola ($\gamma_{0}$).
@@
Interestingly, the model that fits the data of 2023-05-03 also provides an excellent framework for fitting the SED observed in the following days, just by reducing the energy of the radiating particles $\gamma_{max}$, applying a softer spectral index $s$, and a more pronounced spectral curvature $r$. Notably, we observe that the IC component in the later days enters the hard X-ray domain, leading to the observed relative enhancement of the hard X-ray flux. This supports the idea that the flaring event can confidently be interpreted as an outburst of SSC radiation, likely associated with the acceleration of a distribution of charged particles within the jet, which subsequently cooled down through radiative losses. Unfortunately, no radio data were collected during the flaring activity presented in this paper. For completeness, Fig. 6 also shows archival, non-simultaneous data extracted from the SSDC SED Builder and NED archives, covering facilities from radio to gamma-rays, to emphasize the flare process.
@@
The tension observed between the first gamma-ray upper limit and the model of 2023-05-08 can be attributed to limited statistics, which may lead to an inappropriate representation of the spectral index in the corresponding bin. Additionally, we need to take into account the not strictly simultaneous nature of the data, since our observations in the optical, UV, and X-ray bands are representative of less than 111\,1hr of exposure each, while the gamma-ray data represent source visibility integrated over $1$ day. For the last daily bin on May 9, despite marginal agreement between the model predictions and the observations, it becomes more challenging to achieve a satisfactory fit within the framework of our simple energy-loss single-zone SSC model. This may be due to the flaring event losing power and no longer dominating over other emission regions within the source, making the single-zone emission assumption no longer suitable to describe the status of the source at later times of the event.

### v2

@@ 3 Neil Gehrels Swift XRT and UVOT data
Following the powerful gamma-ray outburst of 3C 216, detected on 2023-05-01 (Giacchino et al., 2023), a Target of Opportunity (ToO) request was sent to the Swift Gamma-Ray Burst Mission (Gehrels et al., 2004). Swift executed five visits of the target, on 2023-05-03, 04, 06, 08 and 09, accumulating observations with both the XRT and UVOT instruments. The observations were executed with a roughly regular spacing of $1.5$ days between subsequent visits. During the observations, the XRT instrument worked in photon counting mode, while UVOT performed a sequence of exposures using the $V$, $B$, $U$, $UW1$, $UM2$ and $UW2$ photometric filters.
3C 216 was detected in all the X-ray/UV/optical bands with values of the X-ray flux of the order of $\Phi_{X}\sim(2.3-3.2)\times 10^{-12}\,\mathrm{erg/cm^{2}/s}$, UV magnitudes in the range $14.6$-$15.8$, and optical ones in the range $15.2$-$16.5$. The results of these observations, concerning the UV, optical and X-ray fluxes, are detailed in Table 3, together with the X-ray photon indices and a comparison with a previous set of observations carried out on 2010-10-21.
Table 3: Table of fluxes observed by Swift for UVOT (in six bands) and XRT (from $0.3$ to $10$ keV). Photon index $\Gamma_{X}$ for spectral model in X-ray. The error are $1\sigma$. Fluxes are in units of $10^{-12}\,\mathrm{erg\,cm^{-2}\,s^{-1}}$
. Time 2010-11-21 2023-05-03 2023-05-04 2023-05-06 2023-05-08 2023/05/09 UVOT UVV $0.06\pm 0.01$ $12.87\pm 0.18$ $11.97\pm 0.22$ $9.18\pm 0.17$ $6.00\pm 0.15$ $4.31\pm 0.11$ UVOT UBB $0.08\pm 0.01$ $14.62\pm 0.25$ $14.67\pm 0.34$ $11.00\pm 0.26$ $6.84\pm 0.22$ $4.97\pm 0.16$ UVOT UUU $0.05\pm 0.01$ $14.78\pm 0.23$ $14.12\pm 0.32$ $11.18\pm 0.23$ $6.70\pm 0.22$ $5.27\pm 0.16$ UVOT UM2 $0.01\pm 0.01$ $16.54\pm 0.28$ $15.96\pm 0.27$ $13.29\pm 0.25$ $8.87\pm 0.26$ $6.31\pm 0.16$ UVOT UW1 $0.02\pm 0.01$ $17.46\pm 0.31$ $17.40\pm 0.31$ $13.48\pm 0.29$ $9.45\pm 0.28$ $7.18\pm 0.19$ UVOT UW2 $0.02\pm 0.01$ $16.37\pm 0.39$ $15.66\pm 0.46$ $14.34\pm 0.44$ $9.30\pm 0.40$ $7.21\pm 0.29$ XRT $1.54\pm 0.41$ $3.24\pm 0.42$ $2.36\pm 0.42$ $2.40\pm 0.49$ $2.26\pm 0.61$ $2.84\pm 0.72$ $\Gamma_{X}$ $1.24\pm 0.57$ $1.91\pm 0.11$ $2.03\pm 0.16$ $1.58\pm 0.17$ $1.62\pm 0.23$ $1.29\pm 0.18$
@@ 3.1 Analysis of Swift data
The UVOT data were processed according to a standard UVOT software analysis, using heasoft-6.32.1555https://heasarc.gsfc.nasa.gov/docs/software/heasoft/. At first, we combined the exposures of each filter with uvotimsum, so to obtain one image per visit per filter. Then we extracted the flux and magnitude in each band-pass, together with their associated errors, using the task uvotdetect. 3C 216 was clearly detected in a high state, achieving a statistical significance larger than $20\sigma$ in all the visits and with all the photometric filters. The first Swift observation recorded a $U$-band magnitude of $(14.67\pm 0.02)\,$mag, more than 3 magnitudes brighter than the archival value of 17.8817.8817.88\,17.88mag reported in literature in the same band (Ryle & Sandage, 1964), while subsequent observations traced a steadily decreasing trend. In spite of the long time elapsed since this archival photometric determination, the source has also been observed more recently in ugriz photometry by the Sloan Digital Sky Survey, resulting in even fainter magnitude values in all pass-bands (Adelman-McCarthy et al., 2008), and by Swift itself in 2010, obtaining magnitudes of the order of $19$. We can therefore conclude that the UVOT magnitudes observed during the gamma-ray outburst correspond to a high state of the source. In order to recover the intrinsic state, we corrected the observed fluxes, accounting for the effects of foreground extinction due to the Milky Way’s interstellar medium. Since 3C 216 is located far away from the Galactic Plane, it is subject to a modest reddening effect, having $A_{V}=0.052$ mag, $A_{B}=0.069$ mag and $A_{U}=0.082$ mag (Schlafly & Douglas, 2011). To correct the $UV$ pass-bands, we derived the extinction coefficients for the $UW2$, $UM2$ and $UW1$ according to Yi et al. (2023).
@@
The XRT data were processed using xrtpipeline v3.7.0666https://heasarc.gsfc.nasa.gov/lheasoft/ftools/headas/xrt.html. For each observation, spectral analysis of the source was performed within a $35$ arcsec radius from the target, with background emission subtracted from an annular region between an inner radius $r_{in}\sim 100$ arcsec and an outer radius $r_{out}\sim 400$ arcsec, centered on the position of 3C 216. The spectra were then binned to ensure a minimum of one count per bin and modeled using the XSPEC v12.13.1e (Arnaud et al., 1999) package, employing the Cash statistic for minimization Cash (1979). Fluxes were extracted in the soft ($[0.5-2]$ keV) and hard ($[2-10]$ keV) bands. The adopted model is a power-law modified by neutral Galactic absorption ($TB_{abs}\times z_{pow}$).
@@
Figure 5: Light curves from UVOT data across six filters (top panel), XRT data in two energy ranges (medium panel), and Fermi-LAT daily binned data (bottom panel) collected during the Swift follow-up campaign
In Fig. 5 we present the light curves across different frequency bands for the period covered by the Swift observations: the UVOT data for the six filters are shown in the top panel, the XRT data for the soft and hard bands in the middle panel, and the LAT data in the bottom panel. The observed trend indicates a stable decreasing evolution in the light curves across all spectral bands, with the exception of the hard X-ray domain, which appears to be roughly constant or to exhibit a slight increase in activity. This behavior is consistent with the expectations of coherent emission from a single zone model. In this case, indeed, we would expect the energetic particles present in the jet to produce initially strong synchrotron radiation up to the UV and soft X-ray domain, while effectively scattering low energy photons up to the $\gamma$-ray domain through inverse Compton processes. However, as time passes and the radiating particles lose energy, the synchrotron radiation would gradually fade off and the inverse Compton scattering would rather produce hard X-ray radiation, instead of $\gamma$-rays.
@@
Figure 6: SED for the broadband emission of 3C 216, spanning from UV to gamma-ray, based on five datasets observed during the flare peak period. The blue line represents the SSC model emission computed with JetSeT. The green dots indicate the data collected by UVOT, XRT, and Fermi-LAT, along with their respective uncertainties. Upper limits are reported when $TS\leq 10$. Light gray points in the background represent archival, non-simultaneous data from literature, specifically from the following sources: 3C, NRAO, Ohio (Dixon, 1970), FIRST at VLA, Green Bank GB6 and North surveys, combined NRAO and Parkes survey, NVSS, VLBA, Planck ERCSC, PCCS1F and PCCS2F catalogs, WMAP catalog, TWOMASS and WISE infrared catalogs, USNO and SDSS2,6,7, optical surveys, Einstein and ROSAT X-ray source catalogs, Swift XRT SWXRT1 and 1SXPS catalogs (D’Elia et al., 2013; Evans et al., 2014), Chandra ACIS source catalog (Massaro et al., 2015), and the Fermi-LAT DR3 source catalog (Abdollahi et al., 2022) data. These data were extracted from the SSDC SED Builder and NED archives.
4 Discussion
The regular $\gamma$-ray monitoring of 3C 216, shown by Fig. 1 suggests that the source entered a high activity state extending over several months. However, the outburst that resulted in the highest observed daily flux developed in a time scale of approximately 2 days and it is therefore likely to originate from a region of limited size. The presence of multi-frequency follow-up observations allows us to obtain a series of snap-shots of the status of the SED in the days immediately after the $\gamma$-ray record state observed on 2023-05-01 and, subsequently, to attempt an interpretation of its origin.
To model the radiative emission, we used JetSet (Tramacere, 2020), an open source C/Python tool for reproducing the radiative and acceleration processes acting in relativistic jets and Galactic objects (beamed and unbeamed), allowing numerical models to be fit to the observed data. JetSet provides models to calculate SEDs obtained by combining synchrotron-self Compton and external Compton (EC), using different types of source seed radiation field, as well as the Cosmic Microwave Background (CMB). We decided to analyze the time variability of the multi-frequency SED to understand the physical process of the flare activity of 3C 216. We chose to construct our base-line model with the data set corresponding to 2023-05-03, because it is the earliest time for which we obtained simultaneous multi-frequency coverage of the source. In principle, there are many possibilities that could explain the observed $\gamma$-ray emission, including hadronic and leptonic processes, contributions from external radiation fields and multi-zone emission scenarios. However, such models are characterized by a rather large number of free parameters and they require rather complicated fine tuning to predict a coherent SED evolution, particularly in the case of multi-zone scenarios. If, on the other hand, a single-zone model is to be preferred, it is very likely that the $\gamma$-ray production mechanism is IC scattering by charged relativistic particles, which will also produce strong synchrotron emission. We therefore decided to reproduce the first available multi-frequency SED in the framework of a SSC model (Jones et al., 1974), testing whether the SEDs observed at later times could be interpreted by the same model, after accounting for energy loss, as it would be expected in this case. Another advantage of the SSC scenario over other possible alternatives is the limited number of model parameters, which can be reduced to $10$ in the case of a log-parabola particle energy distribution.
Table 4: Geometric and magnetic properties of the emitting region (fixed model parameters).
@@ 4 Discussion
Our strategy consisted in dividing the model parameter in a set of $5$ fixed values, listed in Table 4 and used to describe the size of the radiating zone, its magnetic field and its bulk motion, and another set of variable values, listed in Table 5, which account for the energy evolution of the radiating particles as a function of time. We found that the observed spectral shape of photons can be reproduced by a charged lepton population with a flat low-energy spectral index and a log-parabola energy distribution $f(\gamma)$. The required density, $N=1000\,\mathrm{cm^{-3}}$, is typical of the immediate environment surrounding an AGN, and the blob size is consistent with a variability timescale on the order of $1$ day in the observed frame. The number of emitting particles per unit volume $N$ is given by:
@@
where we adopted a constant value of $\gamma_{min}=2$ for the minimum Lorentz factor of the emitting particles energy distribution. The spectral distribution of particles is described by a log-parabola:
@@
Table 5: List of JetSet model parameter values that evolve over time. The reported parameters include the time, the maximum particle Lorentz factor for the energy distribution of the emitting particles ($\gamma_{\mathrm{max}}$), the curvature of log-parabola ($r$), the spectral index of log-parabola ($s$), and the reference energy of log-parabola ($\gamma_{0}$). The minimum particle Lorentz factor $\gamma_{\mathrm{min}}$ is assumed to be always $2.0$.
Obs. Date $\gamma_{\mathrm{min}}$ $\gamma_{\mathrm{max}}$ $r$ $s$ $\gamma_{0}$
2023-05-03 2.0 $5\times 10^{5}$ $1.05$ $1$ 580
2023-05-04 2.0 $4\times 10^{5}$ $1.05$ $1.05$ $575$
2023-05-06 2.0 $4\times 10^{5}$ $1.07$ $1.07$ $540$
2023-05-08 2.0 $3\times 10^{5}$ $1.15$ $1.15$ $540$
2023-05-09 2.0 $3\times 10^{5}$ $1.15$ $1.2$ $530$
@@
A comparison of the adopted model with JetSeT (Massaro et al., 2006; Tramacere et al., 2009, 2011) and the broadband observed SED is illustrated in Fig. 6. Interestingly, the model that fits the data of 2023-05-03 also provides an excellent baseline for fitting the SED observed in the following days, except for some tension on the $\gamma$-ray emission seen on 2023-05-08, just by reducing the energy of the radiating particles $\gamma_{max}$, applying a softer spectral index $s$, and a more pronounced spectral curvature $r$. Notably, we observe that the IC component in the later days enters the hard X-ray domain, in agreement with the non decreasing trend seen only for this band in Fig. 5. This supports the idea that the peak of the flare can be interpreted as a single-zone outburst, where SSC radiation played a dominant role, likely due to the acceleration of a distribution of charged particles within the jet, which subsequently cooled down through radiative losses. Unfortunately, no radio data were collected during the flaring activity presented in this paper, therefore we can neither verify nor exclude that other processes were involved, at least on the long term. For completeness, Fig. 6 also shows archival, non-simultaneous data extracted from the SSDC SED Builder and NED archives, covering facilities from radio to gamma-rays, to emphasize the magnitude of the peak activity with respect to the historical properties of 3C 216.
@@
The tension observed between the first gamma-ray upper limit and the model of 2023-05-08 can be attributed to limited statistics, which may lead to an inappropriate representation of the spectral index in the estimate of the lowest energy $\gamma$-ray UL. Additionally, we need to take into account the non strictly simultaneous nature of the data, since our observations in the optical, UV, and X-ray bands are representative of less than 111\,1hr of exposure each, while the gamma-ray data represent the source visibility integrated over a full day. For the last daily bin on May 9, despite marginal agreement between the model predictions and the observations, it becomes more challenging to achieve a satisfactory fit within the framework of our simple energy-loss single-zone SSC model. This may be due to the flaring event losing power and being no longer dominant over other emission regions within the source, thus leading to a break-down of the single-zone emission assumption at later times of the event.